OpenStudy (anonymous):

integrate ((3x-1)/(x^2-x-1))

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 3x-1 }{ x^2-x-1}\]

OpenStudy (amistre64):

what is the derivative of the bottom?

OpenStudy (anonymous):

2x-1

OpenStudy (amistre64):

hmmm, then my first idea wont work

OpenStudy (amistre64):

might have to go ahead and partialdecomp it

OpenStudy (amistre64):

what is the factorization of the denom?

OpenStudy (anonymous):

(x-1/2)^2 -3/4

OpenStudy (amistre64):

..... you are not gonna like what the wolf brought up

OpenStudy (anonymous):

i been there couldn't make head or tail of it

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+%283x-1%29%2F%28x%5E2-x-1%29 this tells me that its goona best be down by decomposition

OpenStudy (amistre64):

x^2-x-1 (x-1/4)^2 -5/4 x = 1/4 + sqrt(5)/4 x = 1/4 - sqrt(5)/4 (x -1/4 -sqrt(5)/4) (x -1/4 + sqrt(5)/4) oy that hurts \[3x-1 =A(x -\frac14 -\frac{\sqrt{5}}4)+B(x -\frac14 + \frac{\sqrt{5}}4)\]

OpenStudy (amistre64):

when x=1/4 + sqrt(5)/4 \[3(\frac14 + \frac{\sqrt{5}}4)-1 =A(0)+B(\frac{2\sqrt{5}}4)\] \[\frac{2(\frac34 + \frac{3\sqrt{5}}4-1)}{\sqrt5} =B\]

OpenStudy (amistre64):

same concept for finding A, then it logs up just fine; but oy what a mess

OpenStudy (amistre64):

did i do the factoring right?

OpenStudy (anonymous):

isn't it (x-1/2)^2-5/4

OpenStudy (amistre64):

yeah, i got 4s underneath and they need to be 2s :/

OpenStudy (amistre64):

same mes tho :)

OpenStudy (anonymous):

yeah, thx a million

OpenStudy (amistre64):

good luck ;)